The following data represents the number of minutes customers at a fast food restaurant had to wait in the drive thru line be.
Answer #1

The for the given data

sample mean, T = 6

Sample standard deviation, s= 4.69

a) Z score for 12 = \frac{x-\bar{x}}{s} = \frac{12-6}{4.69} = 1.28

b)Z score for 3 = \frac{x-\bar{x}}{s} =  \frac{3-6}{4.69} = -0.64

c) Z score for 2 = \frac{x-\bar{x}}{s} = \frac{2-6}{4.69} = -0.85

d)Z score for 6 = \frac{x-\bar{x}}{s} = \frac{6-6}{4.69} = 0